# Some stability results for the fractional differential equations with two delays

### v1

@@ Abstract
This paper investigates the stability properties of a nonlinear fractional differential equation with two discrete delays and a delay-dependent coefficient. Such equations arise in various biological and control systems where temporal delays influence feedback mechanisms. Two cases are analyzed: one where the first delay is zero, and the second acts through an exponential coefficient, and another where the first delay is fixed. We derive delay-independent stability conditions using linearization, characteristic equations, and bifurcation theory, along with complete theoretical proofs. The results are validated through numerical simulations and stability diagrams.
@@ 1 Introduction
In 1695, Leibnitz proposed a notation $\frac{d^{n}x}{dt^{n}}$ for the $n^{th}$ order derivative of a function $x(t)$, where $n$ is a natural number. In his letter dated 30 September 1695, L’Hopital leibniz1860leibnizens ; podlubny1998fractional asked the meaning of this term when $n=1/2$. This is popularly known as the birth of the fractional calculus, a branch of mathematics that studies operators of noninteger order. By Fractional Derivative (FD), one means the derivative of an arbitrary order, e.g, a real or complex number, a function of t, etc.. The fractional derivative operators are non-local, unlike the classic integer-order derivative bhalekar2019can ; li2007remarks . Thus, one must provide the state values from the initial point when evaluating the FDs. This procedure proved extremely useful while modeling the memory properties in natural systems Podlubny1999 ; kilbas2006theory . This also gives the flexibility of selecting an order that fits the experimental data mainardi2022fractional ; bagley1983theoretical .
@@
Similarly, the Delay Differential Equations (DDE) contain the states at the past time values smith2010introduction ; lakshmanan2011dynamics . Though the DDEs look very similar to their non-delayed counterparts, they are extremely difficult to analyze analytically hale2013introduction . A single scalar DDE can have oscillatory solutions. Moreover, one can have a scalar DDE exhibiting chaos sprott2007simple ; senthilkumar2005bifurcations ; ruiz2013chaos . The DDEs are infinite-dimensional dynamical systems that occur in various models in physics otto2019nonlinear ; bazighifan2021oscillation , chemistry epstein1991differential ; roussel1996use , biology and medicine rihan2021delay ; bani2017analysis ; glass2021nonlinear ; dong2023application , engineering and control systems insperger2011semi ; michiels2010control ; gumussoy2014computer , finance nagarajan2025stock ; flore2019feynman ; agrawal2020jump , and so on.
@@
The fractional order differential equations with delay(FDDE) contain both these tools: the fractional derivative as well as the delay. Thus, one gets a better model for the given natural system. The numerical methods to solve FDDEs are presented in bhalekar2011predictor ; daftardar2014new . The stability and bifurcation analysis of these systems is developed in bhalekar2016stability ; bhalekar2011fractional ; gupta2024fractional ; bhalekar2022stability . The examples of chaotic FDDEs are bhalekar2012dynamical ; daftardar2012dynamics ; bhalekar2012generalized . The DDEs involving multiple delays occur in applications involving two or more processes gu2003stability ; belair1995age ; niculescu2002delay . The logistic model with two delays is studied in braddock1983two . Piotrowska proposed a model with two delays in tumor growth piotrowska2008hopf . Boullu et al employed such equations in the production of platelets boullu2020stability .
@@
There are very few papers devoted to the analysis of such equations. Hale and Huang hale1993global , Belair and Campbell belair1994stability , Li, Ruan, and Weili1999stability , Bhalekar bhalekar2019analysing studied some particular cases of DDEs with multiple delays. Bhalekar and Dutta bhalekar2025analysis discussed the stability of a class of FDDEs with two delays. To the best of our knowledge, such equations with fractional derivatives have not been studied much in the literature. This motivates us to work on their stability and bifurcation analysis. In this work, we provide some delay-independent results.
@@
This paper is organized as follows. Section 2 reviews preliminary results on fractional delay differential equations with a single discrete delay. In Section 3, we introduce the nonlinear fractional delay differential equation with delay-dependent coefficients, establish its equilibrium, and derive the corresponding linearized and characteristic equations. Sections 4 and 5 present the main stability theorems for the cases $\tau_{1}=0$ and $\tau_{1}=1$, together with illustrative examples and graphical results. Finally, Section 6 provides concluding remarks and outlines possible directions for future research.
@@ 2 Preliminaries
We consider a linear delay differential equation with Caputo fractional derivative $D^{\alpha}$ podlubny1998fractional ; kilbas2006theory .
@@
bhalekar2016stability Consider the scalar FDE with a single discrete delay:
@@ 3 The Main Model
where $D^{\alpha}$ denotes the Caputo fractional derivative, $\gamma\in\mathbb{R}$, and $g\in C^{1}(\mathbb{R})$. The parameters $\tau_{1},\tau_{2}\geq 0$ represent discrete delays, and the exponential factor $e^{-\gamma\tau_{2}}$ introduces a delay-dependent coefficient. The integer-order counterpart of this equation, analyzed in boullu2020stability , represents a model for platelet production.
@@
We investigate the stability properties of the linearized system (3) in the following particular cases:
  • Case 1: $\tau_{1}=0$ and arbitrary $\tau_{2}>0.$
  • Case 2: $\tau_{1}=1$ and arbitrary $\tau_{2}>0.$
@@ 4 Case 1: $\tau_{1}=0$
The above equation is of the general form
@@ 4.1 Illustrative Examples for Case 1
5 Case 2: $\tau_{1}=1$
In this case, both delays are retained. Substituting $\tau_{1}=1$ and $\tau_{2}=\tau$ in eq.(3), we obtain
$D^{\alpha}x(t)=-\gamma x(t)+kx(t-1)-ke^{-\gamma\tau}x(t-\tau-1),\quad 0<\alpha\leq 1.$ (7)
The corresponding characteristic equation is given by
$\lambda^{\alpha}=-\gamma+ke^{-\lambda}-ke^{-\gamma\tau}e^{-(\tau+1)\lambda}.$ (8)
@@ 5 Case 2: $\tau_{1}=1$
Consider equation (7) with parameters $k,\gamma\in\mathbb{R}$ and $\tau\geq 0$. Then the equilibrium point $x_{*}=0$ of eq.(2) exhibits the following stability behaviors (see Figure 15):
@@
    Delay-Independent Stability: The equilibrium is asymptotically stable for all $\tau\geq 0$ if either
@@
      If $\gamma<0$ and $k<0$, i.e. $(k,\gamma)$ lies in the third quadrant of the $k-\gamma$ plane, then the equilibrium is unstable for all $\tau\geq 0$.
@@
      For $\alpha=1,$ the equilibrium is unstable in the fourth quadrant.
@@
(a) Consider first $\tau=0$. The eq.(7) reduces to:
@@
with characteristic equation $\lambda^{\alpha}=-\gamma$. If $\gamma>0$, then $\gamma<0$, so $\arg(-\gamma)=\pi>\frac{\alpha\pi}{2}$, and the equilibrium is asymptotically stable at $\tau=0$.
Now, assume $\tau>0$, and stability will change if a root of the characteristic equation crosses the imaginary axis. Suppose a purely imaginary root $\lambda=iv$ of eq.(8) exists with $v>0$. Substituting into the characteristic equation, we get
$(iv)^{\alpha}=-\gamma+ke^{-iv}\left(1-e^{-(iv+\gamma)\tau}\right).$
@@
$v^{\alpha}\cos\left(\frac{\alpha\pi}{2}\right)$ $=-\gamma+k\cos v-ke^{-\gamma\tau}\cos(v+v\tau),$ (9)
$v^{\alpha}\sin\left(\frac{\alpha\pi}{2}\right)$ $=-k\sin v+ke^{-\gamma\tau}\sin(v+v\tau).$ (10)
For the system to be unstable, there must exist a root satisfying both eq.(9) and (10).
Case (i) Suppose $\gamma>2k>0,\;k>0$. Then using bounds $\cos v\leq 1$ and $\cos(v+v\tau)\leq 1$, it can be estimated that
$k\cos v-ke^{-\gamma\tau}\cos(v+v\tau)\leq k+ke^{-\gamma\tau},$
@@
$\text{RHS}\leq-\gamma+k+ke^{-\gamma\tau}.$
Since $\tau>0$ and $\gamma>0$, we know that $e^{-\gamma\tau}<1$. Therefore,
$\gamma+k+ke^{-\gamma\tau}<-\gamma+k+k=-\gamma+2k.$
@@
$\gamma+k+ke^{-\gamma\tau}<-\gamma+2k<0,$
@@
Case (ii) Here, $\gamma>-2k>0,\;k<0$. A similar argument as in case (i) will be used here. The bounds $\cos v\geq-1$ and $\cos(v+v\tau)\leq 1$ implies
$k\cos v-ke^{-\gamma\tau}\cos(v+v\tau)\leq-k-ke^{-\gamma\tau}.$
@@
$\text{RHS}\leq-\gamma-k-ke^{-\gamma\tau}.$
@@
$\gamma-k-ke^{-\gamma\tau}<-\gamma-2k<0.$
Because the LHS of eq.(9) is non-negative. This gives a contradiction, so no imaginary root exists, and the system will remain stable in this region too.
Case (i) Assume $\gamma<0$ and $k<0$. At $\tau=0$, we have $\lambda^{\alpha}=-\gamma>0$, so $\lambda$ has a positive real root, implying instability. Now for $\tau>0$, define
$F(\lambda):=\lambda^{\alpha}+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda},\text{the characteristic function.}$ (11)
@@
$F(0)=\gamma-k+ke^{-\gamma\tau}.$
$\because\;\;\gamma<0,\tau>0$,
$\qquad\;\;e^{-\gamma\tau}>1$
$\implies ke^{-\gamma\tau}<k\qquad(\because k<0)$
$\implies ke^{-\gamma\tau}-k<0$
$\implies\gamma-k+ke^{-\gamma\tau}<\gamma<0$
@@
$F(\lambda):=\lambda+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda}.$
Now, for the fourth quadrant, $\gamma<0$, hence, $\gamma>0$. Also,
$F(-\gamma):=0.$
Since the characteristic function has a positive real root, hence equilibrium is unstable in this case. ∎
@@
In the fourth quadrant of the $k\!-\!\gamma$ plane (so $k>0$ and $\gamma<0$), fix the parameters $\alpha>0$ and $\tau>0$. There exists a critical value
$k_{*}\;=\;\frac{(1+\tau)^{1+\tfrac{1}{\tau}}}{\tau e^{\gamma}}\left[\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma\right],$ (12)
@@
$F(\lambda)=\lambda^{\alpha}+\gamma-ke^{-\lambda}+ke^{-(\lambda+\gamma)\tau-\lambda}$
@@
where $\quad L(\lambda)=\lambda^{\alpha}+\gamma\quad\text{and}\quad R(\lambda)=ke^{-\lambda}\big(1-e^{-(\lambda+\gamma)\tau}\big).\\$ Differentiating $R(\lambda)$ gives
$R^{\prime}(\lambda)=ke^{-\lambda}\Big((1+\tau)e^{-(\lambda+\gamma)\tau}-1\Big).$
@@
$\lambda_{*}=\frac{\ln(1+\tau)}{\tau}-\gamma>0.$
@@
$R^{\prime\prime}(\lambda)=ke^{-\lambda}(1-e^{(-\gamma-\lambda)\tau}(1+\tau)^{2}).$
@@
$R^{\prime\prime}(\lambda_{*})=-k\,\tau\,e^{\gamma}(1+\tau)^{-1/\tau}<0.$
@@
$\therefore\qquad\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma=\frac{k_{*}\tau e^{\gamma}}{(1+\tau)^{\frac{1}{\tau}+1}}$
$\implies k_{*}=\frac{(1+\tau)^{1+\tfrac{1}{\tau}}}{\tau e^{\gamma}}\left[\left(\frac{\ln(1+\tau)}{\tau}-\gamma\right)^{\alpha}+\gamma\right].$
Figure 14: Comparison of $L(\lambda)=\lambda^{\alpha}+\gamma$ and $R(\lambda)$. At $k=k_{*}$ they intersect at $\lambda_{*},$ where $R(\lambda)$ attains its maximum. For $k>k_{*},$ $R(\lambda)$ exceeds $L(\lambda)$ near $\lambda_{*}$, ensuring a positive root of $F(\lambda)$ and instability.
Finally, since $R(\lambda)$ has a unique maximum, if $k>k_{*}$ then $\max_{\lambda>0}R(\lambda)=R(\lambda_{*})=\frac{k\tau e^{\gamma}}{(1+\tau)^{\frac{1}{\tau}+1}}>L(\lambda_{*})$. Thus $F(\lambda)=L(\lambda)-R(\lambda)$ becomes negative near $\lambda_{*}$, but for large $\lambda$ one has $F(\lambda)\sim\lambda^{\alpha}>0$. By continuity, there exists some $\lambda>0$ with $F(\lambda)=0$, i.e. a positive root. This root implies instability of the equilibrium whenever $k>k_{*}$. ∎
Figure 15: Stability regions for $\tau_{1}=1$

### v2

@@ Abstract
This paper investigates the stability properties of a nonlinear fractional differential equation with two discrete delays and a delay-dependent coefficient. Such equations arise in various biological and control systems where temporal delays influence feedback mechanisms. In the first case, we set one of the delays to zero and analyzed the resulting system. We then extended the study to a more general case where both delays are allowed to vary. We derive delay-independent stability conditions using linearization, characteristic equations, and bifurcation theory, along with complete theoretical proofs. The results are validated through numerical simulations and stability diagrams.
Keywords: Fractional Calculus; stability; delay.
@@ 1 Introduction
In 1695, Leibnitz proposed a notation $\frac{d^{n}x}{dt^{n}}$ for the $n^{th}$ order derivative of a function $x(t)$, where $n$ is a natural number. In his letter dated 30 September 1695, L’Hopital Leibniz (1860); Podlubny (1998) asked the meaning of this term when $n=1/2$. This is popularly known as the birth of the fractional calculus, a branch of mathematics that studies operators of noninteger order. By Fractional Derivative (FD), one means the derivative of an arbitrary order, e.g, a real or complex number, a function of t, etc.. The fractional derivative operators are non-local, unlike the classic integer-order derivative Bhalekar and Patil (2019); Li and Deng (2007). Thus, one must provide the state values from the initial point when evaluating the FDs. This procedure proved extremely useful while modeling the memory properties in natural systems Podlubny (1999); Kilbas et al. (2006). This also gives the flexibility of selecting an order that fits the experimental data Mainardi (2022); Bagley and Torvik (1983).
@@
Similarly, the Delay Differential Equations (DDE) contain the states at the past time values Smith (2010); Lakshmanan and Senthilkumar (2011). Though the DDEs look very similar to their non-delayed counterparts, they are extremely difficult to analyze analytically Hale and Lunel (2013). A single scalar DDE can have oscillatory solutions. Moreover, one can have a scalar DDE exhibiting chaos Sprott (2007); Senthilkumar and Lakshmanan (2005); Ruiz-Herrera (2013). The DDEs are infinite-dimensional dynamical systems that occur in various models in physics Otto et al. (2019); Bazighifan and Askar (2021), chemistry Epstein and Luo (1991); Roussel (1996), biology and medicine Rihan (2021); Bani-Yaghoub (2017); Glass et al. (2021); Dong et al. (2023), engineering and control systems Insperger and Stépán (2011); Michiels et al. (2010); Gumussoy and Gahinet (2014), finance Nagarajan et al. (2025); Flore and Nappo (2019); Agrawal and Hu (2020), and so on.
@@
The fractional order differential equations with delay(FDDE) contain both these tools: the fractional derivative as well as the delay. Thus, one gets a better model for the given natural system. The numerical methods to solve FDDEs are presented in Bhalekar and Daftardar-Gejji (2011); Daftardar-Gejji et al. (2014). The stability and bifurcation analysis of these systems is developed in Bhalekar (2016); Bhalekar et al. (2011); Gupta and Bhalekar (2024); Bhalekar and Gupta (2022). The examples of chaotic FDDEs are Bhalekar (2012); Daftardar-Gejji et al. (2012); Bhalekar et al. (2012). The DDEs involving multiple delays occur in applications involving two or more processes Gu et al. (2003); Bélair et al. (1995); Niculescu (2002). The logistic model with two delays is studied in Braddock and Van den Driessche (1983). Piotrowska proposed a model with two delays in tumor growth Piotrowska (2008). Boullu et al employed such equations in the production of platelets Boullu et al. (2020).
@@
There are very few papers devoted to the analysis of such equations. Hale and Huang Hale and Huang (1993), Belair and Campbell Bélair and Campbell (1994), Li, Ruan, and WeiLi et al. (1999), Bhalekar Bhalekar (2019) studied some particular cases of DDEs with multiple delays. Bhalekar and Dutta Bhalekar and Dutta (2025) discussed the stability of a class of FDDEs with two delays. To the best of our knowledge, such equations with fractional derivatives have not been studied much in the literature. This motivates us to work on their stability and bifurcation analysis. In this work, we provide some delay-independent results.
@@
This paper is organized as follows. Section 2 reviews preliminary results on fractional delay differential equations with a single discrete delay. In Section 3, we introduce the nonlinear fractional delay differential equation with delay-dependent coefficients, establish its equilibrium, and derive the corresponding linearized and characteristic equations. Sections 4 and 5 present the main stability theorems for the cases $\tau_{1}=0$ and $\tau_{1}>0$ and $\tau_{2}\geq 0$ respectively, together with illustrative examples and graphical results. Finally, Section 6 provides concluding remarks and outlines possible directions for future research.
@@ 2 Preliminaries
We consider a linear delay differential equation with Caputo fractional derivative $D^{\alpha}$ Podlubny (1998); Kilbas et al. (2006).
@@
Bhalekar (2016) Consider the scalar FDE with a single discrete delay:
@@ 3 The Main Model
where $D^{\alpha}$ denotes the Caputo fractional derivative, $\gamma\in\mathbb{R}$, and $g\in C^{1}(\mathbb{R})$. The parameters $\tau_{1},\tau_{2}\geq 0$ represent discrete delays, and the exponential factor $e^{-\gamma\tau_{2}}$ introduces a delay-dependent coefficient. The integer-order counterpart of this equation, analyzed in Boullu et al. (2020), represents a model for platelet production.
@@ 4 Case 1: $\tau_{1}=0$
Note that even though the eq. (5) is a scalar DDE with a single delay, it is not same as that of eq. (1) studied in the literature. The coefficient of $x(t-\tau)$ in (5) depends on the delay, unlike (1). The eq. (5) is of the general form
@@ 4.1 Illustrative Examples for Case 1
5 Case 2: $\tau_{1}>0,\;\tau_{2}\geq 0$
[…]
@@ 5 Case 2: $\tau_{1}>0,\;\tau_{2}\geq 0$
The equilibrium point $x_{*}=0$ of eq. (2) exhibits the following stability behaviors (see Fig. 15):
@@
    Delay-Independent Stability: The equilibrium is asymptotically stable for all delay values $\tau_{1}\geq 0,\;\tau_{2}\geq 0$ and for $0<\alpha\leq 1$. if either
@@
      If $\gamma<0$ and $k<0$, i.e. $(k,\gamma)$ lies in the third quadrant of the $k\gamma$-plane, then the equilibrium is unstable for all $\tau_{1}\geq 0,\;\tau_{2}\geq 0$ and for $0<\alpha\leq 1$.
@@
      For $\alpha=1,$ the equilibrium is unstable in the fourth quadrant of the $k\gamma$-plane.
@@
(a) Consider first $\tau_{2}=0$. The eq.(3) reduces to:
@@
with characteristic equation $\lambda^{\alpha}=-\gamma$. If $\gamma>0$, then $\gamma<0$, so $\arg(-\gamma)=\pi>\frac{\alpha\pi}{2}$, and the equilibrium is asymptotically stable at $\tau_{2}=0$.
Now, assume $\tau_{2}>0$, and stability will change if a root of the characteristic equation crosses the imaginary axis. Suppose a purely imaginary root $\lambda=iv$ of eq.(4) exists with $v>0$. Substituting into the characteristic equation, we get
$(iv)^{\alpha}=-\gamma+ke^{-iv\tau_{1}}\left(1-e^{-(iv+\gamma)\tau_{2}}\right).$
@@
$v^{\alpha}\cos\left(\frac{\alpha\pi}{2}\right)$ $=-\gamma+k\cos v\tau_{1}-ke^{-\gamma\tau_{2}}\cos(v(\tau_{1}+\tau_{2})),$ (7)
$v^{\alpha}\sin\left(\frac{\alpha\pi}{2}\right)$ $=-k\sin v\tau_{1}+ke^{-\gamma\tau_{2}}\sin(v(\tau_{1}+\tau_{2})).$ (8)
For the system to be unstable, there must exist a root $v>0$ satisfying both eq.(7) and (8).
Case (i) Suppose $\gamma>2k>0,\;k>0$. Then using bounds $\cos v\tau_{1}\leq 1$ and $\cos(v(\tau_{1}+\tau_{2}))\leq 1$, it can be estimated that
$k\cos v\tau_{1}-ke^{-\gamma\tau_{2}}\cos(v(\tau_{1}+\tau_{2}))\leq k+ke^{-\gamma\tau_{2}},$
@@
$\text{RHS}\leq-\gamma+k+ke^{-\gamma\tau_{2}}.$
Since $\tau_{2}>0$ and $\gamma>0$, we know that $e^{-\gamma\tau_{2}}<1$. Therefore,
$\gamma+k+ke^{-\gamma\tau_{2}}<-\gamma+k+k=-\gamma+2k.$
@@
$\gamma+k+ke^{-\gamma\tau_{2}}<-\gamma+2k<0,$
@@
Case (ii) Here, $\gamma>-2k>0,\;k<0$. A similar argument as in case (i) will be used here. The bounds $\cos v\tau_{1}\geq-1$ and $\cos(v(\tau_{1}+\tau_{2}))\leq 1$ implies
$k\cos v\tau_{1}-ke^{-\gamma\tau_{2}}\cos(v(\tau_{1}+\tau_{2}))\leq-k-ke^{-\gamma\tau_{2}}.$
@@
$\text{RHS}\leq-\gamma-k-ke^{-\gamma\tau_{2}}.$
@@
$\gamma-k-ke^{-\gamma\tau_{2}}<-\gamma-2k<0.$
Because the LHS of eq.(7) is non-negative, this gives a contradiction. So no imaginary root exists, and the system will remain stable in this region too.
Case (i) Assume $\gamma<0$ and $k<0$. At $\tau_{2}=0$, we have $\lambda^{\alpha}=-\gamma>0$, so $\lambda$ has a positive real root, implying instability. Now for $\tau_{2}>0$, define
$F(\lambda):=\lambda^{\alpha}+\gamma-ke^{-\lambda\tau_{1}}+ke^{-(\lambda+\gamma)\tau_{2}-\lambda\tau_{1}},\text{the characteristic function.}$ (9)
@@
$F(0)=\gamma-k+ke^{-\gamma\tau_{2}}.$
$\because\;\;\gamma<0,\tau_{2}>0$,
$\qquad\;\;e^{-\gamma\tau_{2}}>1$
$\implies ke^{-\gamma\tau_{2}}<k\qquad(\because k<0)$
$\implies ke^{-\gamma\tau_{2}}-k<0$
$\implies\gamma-k+ke^{-\gamma\tau_{2}}<\gamma<0$
@@
$F(\lambda):=\lambda+\gamma-ke^{-\lambda\tau_{1}}+ke^{-(\lambda+\gamma)\tau_{2}-\lambda\tau_{1}}.$
Now, for the fourth quadrant, $\gamma<0$, hence, $\gamma>0$. Also, $F(-\gamma)=0$.
Since the characteristic function has a positive real root viz. $\gamma$, hence the equilibrium is unstable in this case. ∎
@@
In the fourth quadrant of the $k\!-\!\gamma$ plane (so $k>0$ and $\gamma<0$), fix the parameters $0<\alpha<1$ and $\tau>0$. There exists a critical value
$k_{*}\;=\frac{(\tau_{1}+\tau_{2})^{1+\frac{\tau_{1}}{\tau_{2}}}}{\tau_{2}\;e^{\gamma\tau_{1}}\;\tau_{1}^{\tau_{1}/\tau_{2}}}\left[\left(\frac{\log\left(1+\frac{\tau_{2}}{\tau_{1}}\right)}{\tau_{2}}-\gamma\right)^{\alpha}+\gamma\right],$ (10)
@@
$F(\lambda)=\lambda^{\alpha}+\gamma-ke^{-\lambda\tau_{1}}+ke^{-(\lambda+\gamma)\tau_{2}-\lambda\tau_{1}}$
@@
where $\quad L(\lambda)=\lambda^{\alpha}+\gamma\quad\text{and}\quad R(\lambda)=ke^{-\lambda\tau_{2}}\big(1-e^{-(\lambda+\gamma)\tau_{1}}\big).\\$ We have $L(0)=\gamma<0$ and $L^{\prime}(\lambda)=\alpha\lambda^{\alpha-1}>0.$ So, $L(\lambda)$ is strictly increasing w.r.t $\lambda.$
Also, $R(0)=k(1-e^{-\gamma\tau_{2}})<0\;\;\because e^{-\gamma\tau_{2}}>1.$ As $\lambda\rightarrow\infty,\;R(\lambda)\rightarrow 0.$
Differentiating $R(\lambda)$ gives
$R^{\prime}(\lambda)=ke^{-\lambda\tau_{1}}\Big((\tau_{1}+\tau_{2})e^{-(\lambda+\gamma)\tau_{2}}-\tau_{1}\Big).$
@@
$\lambda_{*}=\frac{1}{\tau_{2}}\log\left(1+\frac{\tau_{2}}{\tau_{1}}\right)-\gamma>0.$
@@
$R^{\prime\prime}(\lambda)=ke^{-\lambda\tau_{1}}({\tau_{1}}^{2}-e^{(-\gamma-\lambda)\tau_{2}}(\tau_{1}+\tau_{2})^{2}).$
@@
$R^{\prime\prime}(\lambda_{*})=-k\,\tau_{1}\,\tau_{2}\,e^{\gamma\tau_{1}}{\left(\frac{\tau_{1}}{\tau_{1}+\tau_{2}}\right)}^{\tau_{1}/\tau_{2}}<0.$
@@
$\therefore$ $\lambda_{*}^{\alpha}+\gamma=k_{*}e^{-\lambda_{*}\tau_{1}}(1-e^{-(\lambda_{*}+\gamma)\tau_{2}})$ (11)
$\implies$ $\left(\frac{1}{\tau_{2}}\log\left(1+\frac{\tau_{2}}{\tau_{1}}\right)-\gamma\right)^{\alpha}+\gamma=\frac{k_{*}\;\tau_{2}\;\tau_{1}^{\frac{\tau_{1}}{\tau_{2}}}\;e^{\gamma\tau_{1}}}{(\tau_{1}+\tau_{2})^{\frac{\tau_{1}}{\tau_{2}}+1}}$ (12)
$\implies$ $k_{*}=\frac{(\tau_{1}+\tau_{2})^{1+\frac{\tau_{1}}{\tau_{2}}}}{\tau_{2}\;e^{\gamma\tau_{1}}\;\tau_{1}^{\tau_{1}/\tau_{2}}}\left[\left(\frac{\log\left(1+\frac{\tau_{2}}{\tau_{1}}\right)}{\tau_{2}}-\gamma\right)^{\alpha}+\gamma\right].$ (13)
Figure 14: Comparison of $L(\lambda)=\lambda^{\alpha}+\gamma$ and $R(\lambda)$. At $k=k_{*}$ they touch each other at $\lambda_{*},$ where $R(\lambda)$ attains its maximum. For $k>k_{*},$ $R(\lambda)$ exceeds $L(\lambda)$ near $\lambda_{*}$, ensuring a positive root of $F(\lambda)$ and instability.
Finally, since $R(\lambda)$ has a unique maximum, if $k>k_{*}$ then
$\max_{\lambda>0}R(\lambda)=R(\lambda_{*})=\frac{k\;\tau_{2}\;\tau_{1}^{\frac{\tau_{1}}{\tau_{2}}}\;e^{\gamma\tau_{1}}}{(\tau_{1}+\tau_{2})^{\frac{\tau_{1}}{\tau_{2}}+1}}>L(\lambda_{*}).$
Thus $F(\lambda)=L(\lambda)-R(\lambda)$ becomes negative near $\lambda_{*}$, but $F(\lambda)$ approaches infinity as $\lambda$ approaches infinity. By continuity, there exists some $\lambda>0$ with $F(\lambda)=0$, i.e. a positive root. This root implies instability of the equilibrium whenever $k>k_{*}$. ∎
Figure 15: Stability regions for $\tau_{1}>0$